# RLEF: Grounding Code LLMs in Execution Feedback with Reinforcement Learning

### v1

@@ A.3 Supervised Fine-Tuning
We fine-tune models for next-token prediction, computing losses on the last response only (i.e., on responses passing both public and private tests); this produced slightly better models compared to training on all responses. We sweep over learning rates $5e^{-7}$ and $2e^{-6}$, and 2 and 3 epochs with a batch size of 64 and sequence length 8192. A linear warmup is performed over 10 steps, and learning rates are annealed according to a cosine schedule. Weight decay is set to 0.1. Models are evaluated after 200 optimizer steps with AdamW and we select final parameters based on validation set performance.
@@ B.1 Pre-trained Models
Table 3: $1@3$ solve rates for few-shot prompting and supervised fine-tuning (SFT) with pre-trained Llama 3.1 models on CodeContests.
@@ B.2 Feedback from Private Tests
We verify whether our RLEF-trained models can benefit from larger test sets during inference by including feedback from private and generated tests. Specifically, we test each model response against all available test cases and provide execution feedback for up to 8 failed test cases. Comparing 1@3 solve rates (temperature 0.2) with a turn limit of 3, the 8B RLEF model can improve from 17.2 to 18.1 on the valid set, whereas on the test set we see a drop from 16.0 to 14.4. For the 70B RLEF model, validation set performance improves from 34.0 to 36.8 and on the test set we obtain 40.9 compared to 38.0 with feedback limited to public tests.

### v2

@@ A.3 Supervised Fine-Tuning
We fine-tune models for next-token prediction, computing losses on the last response only (i.e., on responses passing both public and private tests); this produced slightly better models compared to training on all responses. We sweep over learning rates $5e^{-6}$ and $2e^{-6}$, and 2 and 3 epochs with a batch size of 64 and sequence length 8192. A linear warmup is performed over 10 steps, and learning rates are annealed according to a cosine schedule. Weight decay is set to 0.1. Models are evaluated after 200 optimizer steps with AdamW and we select final parameters based on validation set performance.
@@ B.1 Pre-trained Models
{NiceTabular}
llrr Method (8B) Valid Test
RLEF 17.2 16.0
No Execution Feedback 12.2 10.9
Token-level Value Function 13.1 13.7
Single-turn RL 10.2 10.9
Single-turn w/ Repair 14.8 12.6
(f) (a) $1$@$3$ solve rates for few-shot prompting and supervised fine-tuning (SFT) with Llama 3.1 Base models on CodeContests. (b) Further results from Llama 3.1 Instruct 8B ($1$@$3$): withholding execution feedback from public training during RL; learning a value function on the token level; training a dedicated code repair model and applying it to outputs of the single-turn RL model.
@@ B.2 Feedback from Private Tests
We verify whether our RLEF-trained models can benefit from larger test sets during inference by including feedback from private and generated tests. Specifically, we test each model response against 20 available test cases, including private tests, and provide execution feedback for up to 8 failed test cases. Comparing $1$@$3$ solve rates (temperature 0.2) with a turn limit of 3, the 8B RLEF model can improve from 17.2 to 18.1 on the valid set, whereas on the test set we see a drop from 16.0 to 14.4. For the 70B RLEF model, validation set performance improves from 37.5 to 40.4, and on the test set we obtain 41.2 compared to 38.0 with feedback limited to public tests.
B.3 Extra Repair Model
Le et al. (2022) implements program repair on top of an RL-trained LLM with two extra models: a “critic” predicts the joint outcome of all unit tests (e.g., success, failure, runtime error) and can be used for ranking and determining promising prefixes, and a “repair” model maps wrong solutions to ground truth solutions. In this spirit, we evaluate the effect of a dedicated repair model to improve the single-turn 8B model from Section 3.4.2 as follows.
During the RL training procedure, we collect all generations that do not pass the public unit tests. For the training duration of 12,000 gradient steps, this amounts to 1.48M samples. Next we construct training dialogues with the original prompt (as described in Section C.1), the wrong generation, and a random correct generation for the respective problem from the CodeContest training set. We apply additional processing to the CodeContest solutions by making sure they do indeed pass the provided unit tests and unifying their indentation. We then train repair models via supervised fine-tuning of Llama 3.1 8B Instruct, sweeping over learning rates $5e^{-6}$, $2e^{-6}$, and $1e^{-6}$, and 1 or 2 epochs with a batch size of 64 and a sequence length of 8192.
For evaluations, we estimate $1$@$3$ solve rates by generating an initial program with the RL-trained model followed by up to two independent samples from the repair model. Similar to our main RLEF setting, we refrain from (further) repair if the latest solution passes the public tests. We evaluate all models from the sweep in intervals of 400 gradient steps and select the best checkpoint based on validation set performance. This checkpoint achieves, in combination with the single-turn RL model, a $1$@$3$ solve rate of 14.8 on the validation set and 12.6 on the test set, which is a significant increase over the single-turn RL model alone (10.2 and 10.9, respectively; Section 3.4) but falls short of the corresponding RLEF-trained model which combines code synthesis and code repair (17.2 and 16.0, respectively; Section 3.2)
B.4 RL Training Without Public Test Execution Feedback
We validate our setup consisting of inline execution feedback and early stopping based on public tests (Section 2.1) with an ablation where we withhold information from public tests. Concretely, we remove execution feedback from the prompt for subsequent solutions (Section C.1), starting directly with “Give it another try”. We always ask the model for two follow-up solutions this way (i.e., for a total of three solutions). We do keep our reward definition from Section 2.2 but do not end episodes when public tests are passing.
The resulting model, starting from Llama 3.1 8B Instruct, obtains a $1$@$3$ solve rate of 12.2 on the validation and 10.9 on the test set (Section B.1). This is better than the initial instruct model (8.9 and 10.2, respectively) but significantly below the corresponding RLEF-trained model (17.2 and 16.0, respectively).
B.5 Token-Level Value Function
Here we do not train a value function the level of responses (Section 2.2, Section A.1) but rather predict a value for each token of a response. Our reward formulation remains unchagned; consequently, due to the discount factor being set to $1$, the value function target (reward-to-go) for each token of a response is the same. However, we now compute separate per-token advantages.
With this approach and otherwise identical settings, we achieve a $1$@$3$ solve rate of 13.1 on the validation and 13.7 on the test set, starting from Llama 3.1 8B Instruct (Section B.1). This is below the 17.2 and 16.0 results with the turn-level value function (Section 3.2).